OpenStudy (anonymous):

Sandra plans to fill a corner of her yard with decorative stones. Each row has six more stones than the prior row. Part 1: If there were five stones in the first row, how many stones were used to fill all 40 rows? Part 2: Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

i think in this one you use arithmetic sequence or series

OpenStudy (anonymous):

an = a1 + (n-1)d or Sn = n/2 (a1+an) or Sn = n/2[2a1 + (n-1)d]

OpenStudy (anonymous):

use the arithmetic series one :D

OpenStudy (anonymous):

Okk can you help me ?

OpenStudy (anonymous):

1 5 2 5+6 3 5+6*2 ... so it looks like 5+6*(r-1) 40 5+6*(40-1) =239 Lets look at the formulas http://www.purplemath.com/modules/series4.htm The formula for the first n terms of an arithmetic sequence, starting with n = 1, is: (n/2)*(a1+an) (40/2)*(5+239)=4880 i'd dbl check tho

OpenStudy (anonymous):

on second thought i am pretty sure i am right :D

OpenStudy (anonymous):

Yes you are (: thanks you so muchh ((:

OpenStudy (anonymous):

ooh another example 1 5 =5 2 5+6 =11 3 5+6*2 =17 ... so it looks like 5+6*(r-1) 40 5+6*(40-1) =239 so the sequence is 5,11,17 ... 40 The series is 5+11+17...+40